OpenStudy (anonymous):

how do i solve. Find the equation of the line through the point (8,-5) that is parallel to 6x+9y=-6

OpenStudy (anonymous):

6x+9y=-6 ===> y = -2/3x-2/3 ===> slope =-2/3 ===> equation would y=-2/3 x +b ===> -5=-16/3+b==> b = 1/3 ===> y =-2/3 x +1/3

OpenStudy (chaise):

Parrelel line: 6x+9y=-6 Let's put this into point slope formula? y=mx+b 6x+6 =-9y 6x/-9 + 6/9 = y y = 2x/-3 + 2/3 Parallel lines have the same slope and there your other line will have a slope of 2x/-9 and have a different y intercept (b value) Your line will pass through (8, -5) -5 = 2x8/-9 + c c = -29/9 The equation to your line will be: y = -2x/3 + -29/9

OpenStudy (anonymous):

let parallel line is 6x+9y=K .. since this line passing through (8,-5) so it will satisfy it.. hence ... 6(8)+9(-5)=k =>48-45=k k=3 hence required parallel line is --> 6x+9y=3

OpenStudy (chaise):

Oops - I made a insignificant mistake in here - forgot to carry a negative. It makes no difference to the answer though. Not sure if you can find it.

OpenStudy (anonymous):

chaise you wrong

OpenStudy (anonymous):

dbsingh thanks